Hello, could you please help me solve this quadratic word problem: One of the earliest known quadratic equations comes from a 3800 year old document from ancient Egypt that now sits in the Berlin Museum in Germany. Translated into modern mathematical notation, the document, now known as the Berlin Papyrus, poses the following set of equations: x^2+y^2=100 y=3/4(x) solve for the values of x and y.
Answer by fractalier(6550)

The is the intersection of a line with a circle...two solutions...for
x^2+y^2=100
and
y=3/4(x)
we can substitute in to the first and get
x^2 + (9/16)x^2 = 100
(25/16)x^2 = 100
x^2 = 64
x = +/-8 and thus y = +/- 6
(8,6) and (-8,-6)